OpenStudy (anonymous):

Find the constant of variation for the quadratic variation. 6y = 9x2 a. 4/9 b. 2/3 c. 3/2 d. -2

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

$$\Huge 6y = 9x^2 $$

OpenStudy (anonymous):

3/2?

OpenStudy (skullpatrol):

A "quadratic" variation has the form $$\Huge y = kx^2 $$

OpenStudy (skullpatrol):

$$\Huge 1y = kx^2 $$

OpenStudy (skullpatrol):

But your question has a numerical coefficient of 6 in front of the y...so ..

OpenStudy (skullpatrol):

$$\Huge 6y = 9x^2 $$ $$\Huge \frac{1}{6}*6y = \frac{1}{6}*9x^2 $$ $$\Huge \frac{1}{\cancel 6}*\cancel 6y = \frac{1}{6}*9x^2 $$ $$\Huge y = \frac{1}{6}*9x^2 $$